OpenStudy (anonymous):

write an equation in slope-intercept form for the line that satisfies the following condition. passes through (12,-4), perpendicular to the graph of y = 7/12x + 22

OpenStudy (anonymous):

slope intercept form is y=mx + b, where m is the slope and b is the y intercept. Solve for the slope and the y intercept. Do know what is the relation between to perpendicular lines' slopes?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

Say we have a line A with slope Q. A line perpendicular to A has a slope of -1/Q (the negative reciprocal). So you know that the line y = 7/12x + 22 is perpendicular to the line you are solving for, so what does that tell you about the slope of the line you are solving for?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Two lines that are perpendicular have slopes that are negative reciprocals of each other, so if you know the line you are looking for is perpendicular to y = 7/12x + 22, what is the slope of the line you are looking for?

OpenStudy (anonymous):

idk im in 10 grade for summer school

OpenStudy (anonymous):

The slope of the perpendicular line is 7/12, right? So what is the slope of the line you are looking for if it is perpendicular to line with a slope of 7/12, using the special relation I told you above.

OpenStudy (anonymous):

i dont get it ill jsut fail.

OpenStudy (anonymous):

I can't straight up tell you the answer; you won't learn anything that way. Why don't you ask your teacher to explain it to you? I am sure he/she will be more that willing to help you and won't fail you.

OpenStudy (anonymous):

we dont have a teacher is summer school we are jsut on the computer all day.

OpenStudy (anonymous):

This is the answer you can see the inverse and negate that nolastudent was talking about. now use the point given to find the constant of the second line \[y=7/12x+22\] \[y=\frac{-12}{7}x-\frac{140}{7}\]

OpenStudy (anonymous):

using \[y=mx+c\] \[-4=\frac{-12 \times 12 } {7} x+c\] \[-4=\frac{-144}{7}+c\]\[0=\frac{-140}{7}+c \] so \[c=\frac{140}{7}\] put that back into the equation to get \[y=\frac{-12}{7}x - \frac{140}{7}\]\[y=\] hope this helps